OpenStudy (anonymous):

The variables x and y vary directly. Write an equation that relates x and y, when x = 4 and y = 2. A. y = -2x B. y = -1/2x C. y = 1/2x D. y = 2x

OpenStudy (saifoo.khan):

\[y = \frac 12x\]

OpenStudy (anonymous):

Direct variation is y = kx where k is some number. 2 = k(4) divide by 4 to find k 1/2 = k put it back into the original y = kx equation y = 1/2x